Tuesday, Mar. 29: OWL Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans


Previously...

  • MCC restore the beam Hall-A
  • Resolved all problems they mentioned in this morning (1st pass RF separator/ Moller Quad power cycle/...)



Specific mission

  • Note:

1/ When Beam is back:
(a) HARP

Hall-A Harp

HALOG: [1]

(b)BPM

 Find out  BPM position (beam center)
Spot++
Procedure is on the white board or Wiki (see bottom link)

New beam info
BPM A (x:y) = (+0.5: -2.1)
BPM E (x:y) = (+0.5: -2.1)

RASTER [2 x 4.4] : (MCC GUI)
RASTER [2x2] (Hall-A)
HALOG: [2]



2/ Any chance of opportunity of good beam for physics
(beam goes straight ! not through chicane)
DVCS: Kin_48_4

- 15uA
- LH2 target
- DVCS ON
- Calo_theta= 10.1(deg), Distance = 2.5m (from target) - LHRS=3.36GeV, angle=24.925(deg), Q1/Q2/Q3=1916.640(readout:1900)/1600/1600 (Q1 regul.OFF)
- RHRS=2.204GeV, angle=48.75(deg)
- PreScaleR : PS1=PS2=PS3=1, && PS8=200
- PreScaleL : PS9 =1, + DVCS config.

Tentative plan:

Tuesday:
Arc energy measurement : Doug (Morning)
Moller measurement : Sasha (Afternoon)

Wednesday:
Establish the beam through Compton chicane: Dave & Yves

Thursday:
MCC beam study



BPM/Spot++ Check

  • After beam lost for considerable time (like beam study, pass change,...)
  • Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++